OpenStudy (anonymous):

write an equation for x=3 and x=-5

OpenStudy (anonymous):

Well, you know that x=3, so x-3 = 0 Likewise: x = -5 so x+5 = 0 Those are your two factors: Then you do (x+5)(x-3) = 0, now just FOIL the (x+5)(x-3) to get the equation you need.